# Strategic Bidding Wars in On-chain Auctions

### 04 Jan 2024 v2

@@ I Introduction
Relays act as trusted facilitators between builders and the proposer, by validating received blocks, selecting the winning block with the highest bid, and forwarding only the winning block header to the proposer when the proposer calls getHeader to request the block [10]. This ensures that proposers cannot steal block content but can commit to proposing the block by signing the respective block header. Unfortunately, the default relay implementation adds significant latency to MEV-boost auctions, has high operating costs, and benefits from colocation. This incentivizes vertical integration which in turn reduces proposer fees and compromises network security [11]. In an effort to address these shortcomings, optimistic relays assume that received blocks are valid in the short-term, and defer actual validation to a later time [12].
@@ II MEV-Boost Auction Model
We consider a set of $N=\{1,\ldots,n\}$ players (or builders) competing in the MEV-Boost auction game. Each player, indexed by $i$, employs a bidding strategy which can be described as a function $\beta_{i}:X\to\mathbb{R}_{+}$ so that the bid of player $i$ at time $t$ is $b_{i,t}=\beta_{i}(x_{t})$, where $x_{t}\in X$ represents a vector of input variables at time $t\geq 0$. These inputs are discussed next, where the terms player and builder are used interchangeably.
@@
  • Latency $\Delta+\Delta_{i}$. The latency causes a possible delay in the relay’s acceptance of bids relative to the players’ bid submissions. We introduce two components that constitute this latency: global delay, influenced by the relay’s processing speed, and individual delay, determined by the player’s network connectivity. Global delay, denoted as $\Delta$, quantifies the delays experienced by the relay in accepting these bids relative to their arrival. Individual delay, denoted as $\Delta_{i}$, represents the delays in the arrival of player $i$’s bids to the relay relative to their submission time.
@@
Bid cancellations and honest proposer: Bid cancellations are achieved by replacing existing bids with lower ones. Honest proposers typically make a single getHeader request at the start of the slot and commit to the block with the highest bid at that moment. However, proposers can call getHeader multiple times, allowing them to initially select a block with a higher bid and maintain this choice even if a subsequent getHeader request reveals a lower bid, rendering the bid cancellation feature ineffective. In our model, we assume proposers are honest, or equivalently that bid cancellation is effective. Builders are aware of the proposer’s honesty and expect the proposer to close the auction at the beginning of the slot.
@@
Strategy Bidding behavior at time $t\leq T$
@@
$b_{i,t}^{\text{last}}=v_{i}(t)_{+}\times\mathbf{1}\{t\geq\theta\}$
@@
$b_{i,t}^{\text{bluff}}=b_{i,\text{bluff}}\times\mathbf{1}\{t<\theta\}+v_{i}(t)_{+}\times\mathbf{1}\{t\geq\theta\}$
@@ IV Strategy effectiveness
Figure 2: Win rate of strategies in Profile 1 (left) and Profile 2 (right) for varying global delays.
@@
Figure 8: Profit per win distribution of adaptive players under varying revealing times of bluff players.

### 17 Mar 2024 v3

@@ I Introduction
Relays act as trusted facilitators between builders and the proposer, by validating received blocks, selecting the winning block with the highest bid, and forwarding only the winning block header to the proposer when the proposer calls getHeader to request the block [10]. This ensures that proposers cannot steal block content but can commit to proposing the block by signing the respective block header. Unfortunately, the default relay implementation adds significant latency to MEV-boost auctions. This incentivizes vertical integration which in turn reduces proposer fees and compromises network security [11]. To address these shortcomings, optimistic relays assume that received blocks are valid in the short-term, and defer actual validation to a later time [12].
@@ II MEV-Boost Auction Model
We consider a set of $N=\{1,\ldots,n\}$ builders competing in the MEV-Boost auction game. Each builder, indexed by $i$, employs a bidding strategy $s$ which can be described as a function $\beta_{s}:X\to\mathbb{R}_{+}$ so that the bid of player $i$ at time $t$ is $b_{i,t}=\beta_{s}(x_{i,t})$, where $x_{i,t}\in X$ represents a vector of input variables at time $t\geq 0$. These inputs are discussed next, where the terms player and builder are used interchangeably.
@@
  • Latency $\Delta+\Delta_{i}$. The latency causes a delay in the relay’s acceptance of bids relative to the players’ bid submissions. We introduce two components that constitute this latency: global delay, influenced by the relay’s processing speed, and individual delay, determined by the player’s network connectivity. Global delay, denoted as $\Delta$, quantifies the delays experienced by the relay in accepting these bids relative to their arrival. Individual delay, denoted as $\Delta_{i}$, represents the delays in the arrival of player $i$’s bids to the relay relative to their submission.
@@
Bid cancellations and honest proposer: Bid cancellations are achieved by replacing existing bids with lower ones. Honest proposers typically make a single getHeader request at the start of the slot and commit to the block with the highest bid at that moment. However, proposers can call getHeader multiple times, allowing them to initially select a block with a higher bid and maintain this choice even if a subsequent getHeader request reveals a lower bid, rendering the bid cancellation ineffective. In our model, we assume proposers are honest, or equivalently that bid cancellation is effective. Builders are aware of the proposer’s honesty and expect the proposer to terminate the auction at the beginning of the slot.
@@
Strategy
Bid value $b_{i,t}$ at time $t\leq T$
@@
$\beta_{\text{last}}(x_{i,t})=v_{i}(t)_{+}\times\mathbf{1}\{t\geq\theta\}$
@@
$\beta_{\text{bluff}}(x_{i,t})=b_{i,\text{bluff}}\times\mathbf{1}\{t<\theta\}+v_{i}(t)_{+}\times\mathbf{1}\{t\geq\theta\}$
@@ IV Strategy effectiveness
Figure 2: Win rate of strategies in Profile 1 (left) and Profile 2 (right) under varying global delays.
@@
Figure 8: Adaptive players’ profitability under revealing times of bluff players.